Andy purchased 120 pencils at the cost of $2.00 per pencil, Hesold 72 of them at the cost of $2.50 per pencil and the remaining at the cost of $2.00 per pencil find his profit or loss percent.
Answer by math_tutor2020(3835)


120 pencils at $2 per pencil will cost a total of 120*2 = 240 dollars.

A = revenue gained from selling 72 pencils at $2.50 per pencil
A = 72*2.50
A = 180

B = revenue gained from selling the remaining 48 at $2 per pencil
B = 48*2
B = 96

total sales = A+B = 180+96 = 276
"sales" is another term for "revenue".

profit = revenue - cost
profit = 276 - 240
profit = 36

profit percentage = (profit/cost)*100%
profit percentage = (36/240)*100%
profit percentage = (0.15)*100%
profit percentage = 15% is the final answer
The result is positive since the profit was positive, i.e. Andy made money rather than lose money overall.
His profit ($36) is 15% of the cost ($240).


Extra info:
If you want the profit margin, then,
profit margin = (profit/sales)*100%
profit margin = (36/276)*100%
profit margin = (0.1304)*100%
profit margin = 13.04% approximately
His profit ($36) is roughly 13.04% of the sales ($276).

Both profit percentage and profit margin help gauge how well a business is performing. They both are almost identical, except for the denominator of the fraction will be "cost" for the first formula, while "sales" is in the denominator for the second formula. Both have profit in the numerator because we're comparing profit against the denominator.